OpenStudy (anonymous):

factor the trinomial completly it is y^3-14y^2+45y

OpenStudy (anonymous):

y^3-14y^2+45y Take out the common factor = y ( y^2 - 14y + 45) Split the middle term = y ( y^2 - 5y - 9y + 45) Take our the common factor again = y [ y(y-5) - 9(y-5)] Again, take out the common factor and group the rest, can you do it?

OpenStudy (anonymous):

wait i m doing

OpenStudy (theviper):

y^3-14y^2+45y Take y as common =>y( y^2 - 14y + 45) Then factor the quadratic equation =>y(y^2-5y-9y+45) =>y[ y(y-5) - 9(y-5)] =>(y)(y-5)(y-9) Understood @shivraj ??